# Scrutinize What We Ignore: Reining In Task Representation Shift Of Context-Based Offline Meta Reinforcement Learning

### v3

@@ 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
In this section, we aim to demonstrate that as the previous optimization framework in Section 4.1 doesn’t model the variation of task representation explicitly, this framework would provide an insufficient condition for monotonic performance improvement. We first show conditions for monotonic performance improvement of this framework.
@@
$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x)-Z^{*}(\cdot|x)|)$ (7)
Proof: See Section 8.2 in the Appendix. ∎
As shown in Corollary 4.4, the monotonic performance improvement can be guaranteed with only better approximation to $Z^{*}(\cdot|x)$. Next, we will introduce the performance difference bound framework to model variation of task representation.
@@
Denote $J^{1}(\theta_{1})$ as the expected return of the policy $\pi(\cdot|s,Z_{1}(\cdot|x);\theta_{1})$ among the expectation of tasks before update of the context encoder and the policy. Similarly, denote $J^{2}(\theta_{2})$ as the expected return of the policy $\pi(\cdot|s,Z_{2}(\cdot|x);\theta_{2})$ among the expectation of tasks after update of the context encoder and the policy. The performance difference bound in the COMRL setting can be defined as
@@
According to Definition 4.5, we need to find the positive $C$ to improve the performance monotonically. To achieve this, we can derive the lower bound of the performance difference.
@@
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]$ (9)
Proof: See Section 8.2 in the Appendix. ∎
For the need of the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (32) be positive. Then, we would have the following condition to achieve monotonicity.
$\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z_{2}(\cdot|x)-Z^{*}(\cdot|x)|+|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|]\geq 0$ (10)
Compared to the condition in Corollary 4.4, Eq. 10 presents an additional term $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ to achieve monotonic performance improvement, which corresponds exactly to the previous ignored impacts stemming from the variation of the task representation. To achieve monotonic performance improvements, the optimization of task representation should consider not only the approximation to $Z^{*}(\cdot|x)$, but also the magnitude of the update. For example, if we assume that $Z_{1}(\cdot|x)$ is trained from scratch and the update process from $Z_{1}(\cdot|x)$ to $Z_{2}(\cdot|x)$ brings $Z_{2}(\cdot|x)$ close to $Z^{*}(\cdot|x)$, then with the condition in Corollary 4.4, the monotonic performance improvement can be easily achieved with small $\epsilon^{*}_{12}$. However, for the condition in Eq. (10), small $\epsilon^{*}_{12}$ may cause the violation of monotonicity as $|Z_{1}(\cdot|x)-Z_{2}(\cdot|x)|$ remains large.
To highlight the impacts of $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$, we name this issue Task Representation Shift. Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement with sufficient policy improvement $\epsilon^{*}_{12}$ on the condition of appropriate encoder updates, as shown in Theorem 4.10.
@@
There exists an $\alpha$ for any given $Z_{1}(\cdot|x)$, $|Z_{2}(\cdot|x)-\tilde{Z}_{2}(\cdot|x)|^{2}\leq\frac{\alpha}{b}$, where $Z_{2}(\cdot|x)$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ randomly sampled from the training dataset based on $Z_{1}(\cdot|x)$ and $\tilde{Z}_{2}(\cdot|x)$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z^{*}(\cdot|x)$ based on $Z_{1}(\cdot|x)$  [^1] We set discrepancy to be subject to an inverse proportional function is because as the training samples increase, $Z_{2}$ and $\tilde{Z}_{2}$ should become similar..
@@
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $vol(\mathbb{Z})$ as the volume of the task representation coverage simplex. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{vol(\mathbb{Z})}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
@@
0:  Offline training datasets $\bm{X}$, initialized policy $\pi_{\theta}$, context encoder $q_{\phi}$ and given task representation shift threshold $\beta$
@@
3:     if Conditions for updating the context encoder are met then
4:        while Accumulated task representation shift is less than $\beta$ do
5:           Sample context from $\bm{X}$
6:           Obtain task representations from $q_{\phi}$ with inputting the sampled context
7:           Compute $\mathcal{L}_{\text{encoder}}$ concerning maximizing $I(Z;M)$
8:           Update $\phi$ to minimize $\mathcal{L}_{\text{encoder}}$
9:        end while
10:     end if
11:     // Update the policy
12:     Detach the task representations
13:     Sample training data from $\bm{X}$
14:     Compute $\mathcal{L}_{\text{policy}}$ via standard offline RL algorithms
15:     Update $\theta$ to minimize $\mathcal{L}_{\text{policy}}$
16:  end for
@@ 4.3 Practical Implementation
To show the impacts of the task representation shift issue, we use different settings to adjust the context encoder. We denote the given batch size as $bs$ and the accumulation steps of task representation shift as $acc$. Our settings include 1) $k=2\times bs,acc=1$, 2) $k=3\times bs,acc=1$, 3) $k=1\times bs,acc=2$, and 4) $k=1\times bs,acc=3$. All settings as well as the original setting, namely $k=1\times bs,acc=1$ are performed for 8 different random seeds.
With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose 3 representative algorithms that are also widely adopted in previous COMRL endeavors. The reconstruction-based algorithm, passing the state, action, and task representation through the decoder to reconstruct the next state and the reward signal, is applied in [zintgraf2019varibad, dorfman2021offline, li2024towards] and is proven to be the lower bound of $I(Z;M)$. The contrastive-based algorithm, using the task representation to compute distance metric learning loss, is applied in [li2020focal, gao2024context, li2024towards] and is proven to be the upper bound of $I(Z;M)$ as well as contain the lower bound of $I(Z;M)$. The cross-entropy-based algorithm is a straightforward discrete approximation to $I(Z;M)$ and is proposed by our work. Specifically, after extracting from the context encoder, the task representation would be passed through a fully connected layer to get the probabilities for each task. Then, the context encoder is trained by the supervision of the task label and back-propagation of the cross-entropy loss. The implementation details of these algorithms are shown in Appendix Section 8.3.
With respect to the policy learning algorithm, to maintain the consistency of previous COMRL works, we directly adopt BRAC [wu2019behavior] to train the policy.
@@ 5.1 Environments Settings
We adopt MuJoCo [todorov2012mujoco] and MetaWorld [yu2020meta] benchmarks to evaluate the algorithms. Following the protocol of UNICORN  [li2024towards], we randomly sample 20 training tasks and 20 testing tasks from the task distribution. We train the SAC [haarnoja2018soft] agent from scratch for each task and use the collected replay buffer as the offline dataset. During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like [li2020focal] that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
@@ 5.2 MAIN RESULTS
Figure 2: Testing returns of different settings to rein in the task representation shift against the original setting concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 illustrates the learning curves of our different settings to rein in the task representation shift against the original setting on 20 testing tasks for each benchmark concerning 3 widely adopted objectives for maximizing $I(Z;M)$. We find that these specific objectives hold unanimous performance demonstration on all benchmarks that the previous training framework, namely $k=1\times bs,acc=1$, cannot provide the best asymptotic performance. These objectives focus solely on the optimization target of the task representation, namely the optimal solution of $I(Z;M)$. While they are guaranteed to produce reasonable results, as shown in Section 4.1, ignoring the impacts of task representation shift still hinders the capability of these algorithms to reach their full potential. As shown by Figure 2, even with minor changes to the algorithms, the performance improvements are substantial.
Figure 3: Testing returns of different manners to rein in the task representation shift against the original manner concerning cross-entropy, contrastive, and reconstruction based objectives for maximizing $I(Z;M)$ on different data qualities of Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
@@ 5.3 Can The Results Show Consistency Across Different Data Qualities?
Figure 3 demonstrates the learning curves of our different settings to rein in the task representation shift against the original setting concerning 3 widely adopted objectives for maximizing $I(Z;M)$ on Ant-Dir. We find that the contrastive-based algorithm fails on the random dataset. Except for this case, the others show consistent results that reining in the task representation shift may lead to better performance improvements.
@@ 6 Discussion
Figure 4: Testing returns of the pre-training scheme against training from scratch of cross-entropy, contrastive, and reconstruction based objectives on Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
@@ 6.1 Can The Pretraining Scheme Be Adopted To Achieve Better Performance Improvement?
To answer this question, we conduct the experiment on Ant-Dir and still use the cross-entropy, contrastive, and reconstruction based objectives to pre-train the context encoder. As shown in Figure 4, there is a significant performance gap between pre-training and training from scratch. To better theoretically interpret this, we introduce the following Corollary.
@@
Denote $Z_{pretrain}(\cdot|x)$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|]\geq 0$ (12)
Proof: See Section 8.2 in the Appendix. ∎
Due to the implementation error and approximation error, $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|$ can not equal to 0. Hence, according to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements. As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $Z^{*}(\cdot|x)$. Notice that there exists the decomposition $|Z_{pretrain}(\cdot|x)-Z^{*}(\cdot|x)|\leq|Z_{pretrain}(\cdot|x)-Z_{mutual}(\cdot|x)|+|Z_{mutual}(\cdot|x)-Z^{*}(\cdot|x)|$. Since the cross-entropy based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
This analysis further enhances the importance of considering the impacts of task representation shift $|Z_{2}(\cdot|x)-Z_{1}(\cdot|x)|$ in the alternating process. Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
@@
Previous COMRL endeavors [li2020focal, yuan2022robust, gao2024context] mostly apply t-SNE on the learned task representation to show the differentiation of the tasks. They hold the belief that better performance corresponds to better differentiation of the tasks. However, based on our analysis, the visualization result only depends on the task representation at convergence and ignores the task representation variation during the whole optimization process. According to the illustration shown in Figure 5, less desirable differentiation results can also lead to better performance. Hence, it is insufficient to imply performance based on such evaluation principles. Nevertheless, we recognize that the visualization result can be seen as an auxiliary metric to assist in determining the task representation.

### v4

@@ 4.2 Monotonic Performance Improvement Concerning Task Representation Shift
In this section, we aim to demonstrate that as the previous optimization framework in Section 4.1 doesn’t model the variation of task representation explicitly, this framework would provide an insufficient condition for monotonic performance improvement. We first show conditions for monotonic performance improvement of the previous framework.
@@
$\epsilon^{*}_{12}\triangleq J^{*}(\theta_{2})-J^{*}(\theta_{1})\geq\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}(|Z(\cdot|x;\phi)-Z(\cdot|x;\phi^{*})|)$ (7)
As shown in Corollary 4.4, $Z(\cdot|x;\phi)$ should be close to $Z(\cdot|x;\phi^{*})$ such that the lower bound is small enough for finding a policy to achieve monotonic performance improvement. Next, we will introduce the performance difference bound framework to model variation of task representation.
@@
For an alternating update process, we denote $J^{1}(\theta_{1})=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi_{1});\theta_{1})}}[R_{m}(s,a)]$ as the expected return of the policy $\pi(\cdot|s,Z(\cdot|x;\phi_{1});\theta_{1})$ before update of the context encoder and the policy. Similarly, denote $J^{2}(\theta_{2})=\mathbb{E}_{m,x}\mathbb{E}_{(s,a)\sim d_{\pi(\cdot|s,Z(\cdot|x;\phi_{2});\theta_{2})}}[R_{m}(s,a)]$ as the expected return of the policy $\pi(\cdot|s,Z(\cdot|x;\phi_{2});\theta_{2})$ after update of the context encoder and the policy. The performance difference bound in the COMRL setting can be defined as
@@
According to Definition 4.5, we need to find a positive $C$ to improve the performance monotonically. To achieve this, we can derive the lower bound of the performance difference.
@@
$J^{2}(\theta_{2})-J^{1}(\theta_{1})\geq\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi^{*})|+|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|]$ (9)
For achieving the monotonic performance improvement guarantee, we further let the right-hand side of Eq. (9) be positive. Then, we would have the following condition to achieve monotonicity.
$\epsilon^{*}_{12}-\frac{2R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[2|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi^{*})|+|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|]\geq 0$ (10)
Compared to the condition in Corollary 4.4, Eq. 10 presents an additional term $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|$ to achieve monotonic performance improvement, which corresponds exactly to the previous ignored impacts stemming from the variation of the task representation. To achieve monotonic performance improvements, the optimization of task representation should consider not only the approximation to $Z(\cdot|x;\phi^{*})$, but also the magnitude of the update. For example, if we assume that $Z(\cdot|x;\phi_{1})$ is trained from scratch and the update process from $Z(\cdot|x;\phi_{1})$ to $Z(\cdot|x;\phi_{2})$ brings $Z(\cdot|x;\phi_{2})$ close to $Z(\cdot|x;\phi^{*})$, then with the condition in Corollary 4.4, the monotonic performance improvement can be easily achieved with small $\epsilon^{*}_{12}$. However, for the condition in Eq. (10), small $\epsilon^{*}_{12}$ may cause the violation of monotonicity as $|Z(\cdot|x;\phi_{1})-Z(\cdot|x;\phi_{2})|$ remains large.
To highlight the impacts of $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|$, we name this issue Task Representation Shift. Under some mild assumptions, we can conclude that it is possible to achieve monotonic performance improvement with sufficient policy improvement $\epsilon^{*}_{12}$ on the condition of appropriate encoder updates, as shown in Theorem 4.10.
@@
There exists an $\alpha$ for any given $Z(\cdot|x;\phi_{1})$, $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\tilde{\phi}_{2})|^{2}\leq\frac{\alpha}{b}$, where $Z(\cdot|x;\phi_{2})$ denotes the context encoder updated by maximizing $I(Z;M)$ with the data size $b$ randomly sampled from the training dataset based on $Z(\cdot|x;\phi_{1})$ and $Z(\cdot|x;\tilde{\phi}_{2})$ denotes the context encoder updated by fitting the empirical distribution on $b$ i.i.d samples from $Z(\cdot|x;\phi^{*})$ based on $Z(\cdot|x;\phi_{1})$  [^2] For more details to justify these assumptions, please refer to Appendix 8.3..
@@
Denote $\kappa$ as $\frac{(1-\gamma)^{2}}{4R_{max}L_{z}}\epsilon^{*}_{12}-\frac{1}{2}\beta$ and $|Z|$ as the cardinality of the task representation space. Given that the context encoder has already been trained by maximizing $I(Z;M)$ to some extent. When meeting Assumption 4.1, 4.7, 4.8 and 4.9, with a probability greater than $1-\xi$, we can get the monotonic performance improvement guarantee by updating the context encoder via maximizing $I(Z;M)$ from at least extra $k$ samples, where:
$k=\frac{1}{\kappa^{2}}(\sqrt{2\log\frac{2^{|Z|}-2}{\xi}}+\sqrt{\alpha})^{2}$ (11)
@@
0:  Offline training datasets $\bm{X}$, initialized policy $\pi_{\theta}$, context encoder $Z_{\phi}$, given task representation shift threshold $\beta$ and given training batch size for context encoder $N_{\text{bs}}$.
@@
3:     Estimate the conditions (e.g.use Eq. (11) to approximate $k$)
4:     if Conditions for updating the context encoder are met (e.g. $k<=N_{\text{bs}}$) then
5:        while Accumulated task representation shift is less than $\beta$ do
6:           Sample context from $\bm{X}$
7:           Obtain task representations from $Z_{\phi}$ with inputting the sampled context
8:           Compute $\mathcal{L}_{\text{encoder}}$ concerning maximizing $I(Z;M)$
9:           Update $\phi$ to minimize $\mathcal{L}_{\text{encoder}}$
10:        end while
11:     end if
12:     // Update the policy
13:     Detach the task representations
14:     Sample training data from $\bm{X}$
15:     Compute $\mathcal{L}_{\text{policy}}$ via standard offline RL algorithms
16:     Update $\theta$ to minimize $\mathcal{L}_{\text{policy}}$
17:  end for
@@ 4.3 Practical Implementation
As outlined in Algorithm 1, the way to rein in the task representation shift can be seen as two aspects, namely 1) when to update the context encoder determined by $k$ and 2) how many times to update the context encoder determined by $\beta$. To cover these two aspects, we introduce two parameters $N_{\text{k}}$ and $N_{\text{acc}}$. Here, $N_{\text{k}}=n$ denotes that the context encoder needs to be updated every $n$ updates of the policy, and $N_{\text{acc}}=n$ denotes when the context encoder needs to be updated, it is updated $n$ times. Our settings include 1) $N_{\text{k}}=2,N_{\text{acc}}=1$, 2) $N_{\text{k}}=3,N_{\text{acc}}=1$, 3) $N_{\text{k}}=1,N_{\text{acc}}=2$, and 4) $N_{\text{k}}=1,N_{\text{acc}}=3$. All settings as well as the original setting, namely $N_{\text{k}}=1,N_{\text{acc}}=1$, are performed for 8 different random seeds.
With respect to the specific context encoder learning algorithms concerning maximizing $I(Z;M)$, we choose three representative algorithms that have been widely adopted in previous COMRL endeavors. The details of these algorithms are shown in Appendix Section 8.4.
Contrastive-based is applied in (Li et al., 2020; Gao et al., 2024; Li et al., 2024) and it is proven to be the upper bound of $I(Z;M)$.
Reconstruction-based is applied in (Zintgraf et al., 2019; Dorfman et al., 2021; Li et al., 2024) and it is proven to be the lower bound of $I(Z;M)$, which is equivalent to CORRO (Yuan & Lu, 2022) under the offline setting.
Cross-entropy-based is proposed in our work and it is a direct approximation w.r.t $I(Z;M)$.
With respect to the policy learning algorithm, to maintain the consistency of previous COMRL works, we directly adopt BRAC (Wu et al., 2019) to train the policy.
@@ 5.1 Environments Settings
We adopt MuJoCo (Todorov et al., 2012) and MetaWorld (Yu et al., 2020) benchmarks to evaluate the algorithms. Following the protocol of UNICORN  (Li et al., 2024), we randomly sample 20 training tasks and 20 testing tasks from the task distribution. We train the SAC (Haarnoja et al., 2018) agent from scratch for each task and use the collected replay buffer as the offline dataset. During the meta-testing phase, we adopt a fully offline setting in a few-shot manner like (Li et al., 2020) that randomly samples a trajectory from the dataset as the context to obtain the task representation and then passes it to the meta-policy to complete the testing in the true environments.
@@ 5.2 MAIN RESULTS
Figure 2: Testing returns of different settings to rein in the task representation shift. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
Figure 2 and Table 3 illustrate the performance of our different settings to rein in the task representation shift on 20 testing tasks. We find that these specific objectives exhibit clear trends across all benchmarks that the best setting showcases statistically significant performance improvements compared to the worst setting, thereby emphasizing the indispensability of considering the task representation shift issue. We also notice that the previous training framework, namely $N_{\text{k}}=1,N_{\text{acc}}=1$ , fails to deliver the best asymptotic performance across all benchmarks. Though $N_{\text{k}}=1,N_{\text{acc}}=1$ can be seen as an implicit way to rein in the task representation shift, the lack of explicit control still hinders the capability of these base algorithms to reach their full potential. Furthermore, as shown in Table 3, we observe that settings with $N_{\text{k}}>1,N_{\text{acc}}=1$ achieve better performance more frequently than those with $N_{\text{k}}=1,N_{\text{acc}}>1$ w.r.t the original setting $N_{\text{k}}=1,N_{\text{acc}}=1$. Based on this observation, we recommend prioritizing the adjustment of $N_{\text{k}}$ when tuning parameters. This not only tends to yield better performance but also offers the advantage of reducing training time. In contrast, increasing $N_{\text{acc}}$ would increase training time due to the need for multiple context encoder updates within a single alternating step.
Figure 3: Testing returns of different settings to rein in the task representation shift on different data qualities in Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
@@ 5.3 Can The Results Show Consistency Across Different Data Qualities?
Figure 3 and Table 3 demonstrate the performance of our different settings to rein in the task representation shift on Ant-Dir. We find that the contrastive-based algorithm fails on the random dataset. Except for this case, the others show consistent results that reining in the task representation shift can lead to better performance improvements.
@@ 6 Discussion
Figure 4: Testing returns of the pre-training scheme against training from scratch in Ant-Dir. Solid curves refer to the mean performance of trials over 8 random seeds, and the shaded areas characterize the standard deviation of these trials.
@@ 6.1 Can The Pretraining Scheme Be Adopted To Achieve Better Performance Improvement?
To answer this question, we conduct the experiment in Ant-Dir and still use the cross-entropy, contrastive, and reconstruction based objectives to pre-train the context encoder. As shown in Figure 4, there is a significant performance gap between pre-training and training from scratch. To better theoretically interpret this, we introduce the following Corollary.
@@
Denote $Z(\cdot|x;\phi^{pretrain})$ as the task representation distribution after pre-training. When meeting Assumption 4.1, the monotonic performance improvement condition for pre-training scheme is:
$\epsilon^{*}_{12}-\frac{4R_{max}L_{z}}{(1-\gamma)^{2}}\mathbb{E}_{m,x}[|Z(\cdot|x;\phi^{pretrain})-Z(\cdot|x;\phi^{*})|]\geq 0$ (12)
While maximizing $I(Z;M)$ can help $Z(\cdot|x;\phi^{pretrain})$ to approach $Z(\cdot|x;\phi^{mutual})$, the sub-optimal gap between $Z(\cdot|x;\phi^{mutual})$ and $Z(\cdot|x;\phi^{*})$ still persists. Furthermore, given the inherent approximation error between $Z(\cdot|x;\phi^{pretrain})$ and $Z(\cdot|x;\phi^{mutual})$, the discrepancy between $Z(\cdot|x;\phi^{pretrain})$ and $Z(\cdot|x;\phi^{*})$ maybe large. Hence, according to Corollary 6.1, pre-training cannot actually achieve monotonic performance improvements. As the task representation cannot be changed under the pre-training scheme, compared to training from scratch, this scheme loses some degrees of freedom to improve performance.
Additionally, this scheme can be intuitively regarded as degrading to the problem that how to design a pre-training algorithm to better approximate $Z(\cdot|x;\phi^{*})$. Notice that there exists the decomposition $|Z(\cdot|x;\phi^{pretrain})-Z(\cdot|x;\phi^{*})|\leq|Z(\cdot|x;\phi^{pretrain})-Z(\cdot|x;\phi^{mutual})|+|Z(\cdot|x;\phi^{mutual})-Z(\cdot|x;\phi^{*})|$. Since the cross-entropy-based objective is the direct approximation of $I(Z;M)$, holding better performance under the pre-training scheme is also in line with expectation.
This analysis further enhances the importance of considering the impacts of task representation shift $|Z(\cdot|x;\phi_{2})-Z(\cdot|x;\phi_{1})|$ in the alternating process. Nevertheless, how to achieve better performance under the pre-training scheme is also an interesting problem.
@@
Previous COMRL endeavors (Li et al., 2020; Yuan & Lu, 2022; Gao et al., 2024) mostly apply t-SNE on the learned task representation to show the differentiation of the tasks. They hold the belief that better performance corresponds to better differentiation of the tasks. However, based on our analysis, the visualization result only depends on the task representation at convergence and ignores the task representation shift during the whole optimization process. According to the illustration shown in Figure 5, less desirable differentiation results can also lead to better performance. Hence, the visualization results may represent the true task distribution but cannot sufficiently imply the final performance.